The sum of the measures of a complement and a supplement of an angle is 200. What is the measure of the angle?
Answer by mananth(16946)

let the angle be x
the complement = 90-x
supplement = 180-x
..
180-x+90-x=200
270-2x=200
-2x=-270+200
-2x=-70
x=35